Electron plasma waves are propagated in a
uniform plasma with KTe = 100 ev, n=1016 m-3
and B=0. If the frequency f is 1.1 GHz, what is
the wavelength in cm?
1
Expert's answer
2020-07-01T15:42:16-0400
"v=\sqrt{\frac{2K}{m}}\\\lambda=\frac{v}{f}=\frac{1}{f}\sqrt{\frac{2K}{m}}\\\lambda=\frac{1}{1.1\cdot10^{9}}\sqrt{\frac{2(100)(1.6\cdot10^{-19})}{9.11\cdot10^{-31}}}=0.00539\ m"